# Universal Predictors for Mixing Time more than Liouvillian Gap

### v1

@@ Appendix A Appendix A: Mixing time analysis for strong bulk dissipation
$H^{0,D}=-i\gamma\sum_{\alpha=1}^{L}(K_{\alpha,L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes K_{\alpha,R}^{*})^{2}.$ (31)
@@
We assume that the Hamiltonian does not commute with the Lindblad operators, which means the perturbation can connect the ground state to excited states. Since the Hamiltonian acts on all sites, the number of excited states that can be connected to any given ground state by the perturbation is proportional to the total system size, $L$.
We denote the matrix elements of the perturbation as $\langle\psi_{0k}|V^{D}|\psi_{1q}\rangle=J$, where $|\psi_{0k}\rangle$ represents a GS with label $k$ and $|\psi_{1q}\rangle$ represents a first excited state with label $q$. Since the diagonal (Hermitian) shift $V^{D}_{0k,0k}$ only affects the real part, it does not affect the Liouvillian gap. The only contribution to the Liouvillian gap is given by the last term in the Eq.(36)
$\Delta\propto-iL\frac{J^{2}}{\gamma}.$
@@ Appendix B Appendix B: Reason for choosing the lowest excited state in the rewriting of mixing time
$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log(\operatorname{Tr}|\sigma_{1}|)-\log{(\frac{2\eta}{|c_{1}|})}\right].$ (37)
@@
A potential concern is that higher excited states, while decaying faster, might have substantially smaller trace norms and could in principle yield a larger estimate of the mixing time. For example, one could imagine that for some mode $(\beta_{2},\sigma_{2})$,
$\beta_{2}^{-1}\log(\mathrm{Tr}|\sigma_{2}|)\;>\;\Delta^{-1}\log(\mathrm{Tr}|\sigma_{1}|),$ (38)
@@
$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|(K_{L}-K_{R}^{*})^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (39)
@@ C.1 Strong-dissipation regime
In the strong-dissipation regime, the unperturbed doubled-space (vectorized) eigenbasis is $\{|\epsilon_{p},\epsilon_{q}\rangle\}\equiv\{|\epsilon_{p}\rangle_{L}\otimes|\epsilon_{q}\rangle_{R}\}$, where $|\epsilon_{s}\rangle$ are simultaneous eigenstates of the (Hermitian) Lindblad operators $\{K_{j}\}$. The leading correction to the state $|\epsilon_{m},\epsilon_{m}\rangle$ induced by the Hamiltonian superoperator
@@
where $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pq}$ denotes the unperturbed doubled-space spectrum.
@@ C.2 Weak-dissipation regime
$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|(K_{L}-K_{R}^{*})^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (54)

### v2

@@ Appendix A Appendix A: Mixing time analysis for strong bulk dissipation
$H^{0,D}=-i\gamma\sum_{\alpha=1}^{L}\left(K_{\alpha,L}\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes K_{\alpha,R}^{*}\right)^{2}.$ (23)
@@
We assume that the Hamiltonian does not commute with the Lindblad operators, which means the perturbation can connect the ground state to excited states. Since the Hamiltonian acts on all sites, the number of excited states that can be connected to any given ground state by the perturbation is proportional to the total system size, $L$. We denote the matrix elements of the perturbation as $\langle\psi_{0k}|V^{D}|\psi_{1q}\rangle=J$, where $|\psi_{0k}\rangle$ represents a GS with label $k$ and $|\psi_{1q}\rangle$ represents a first excited state with label $q$.
Since the smallest diagonal shift $V^{D}_{0k,0k}$ vanishes, the Liouvillian gap is not determined by the first-order (diagonal) correction from $V^{D}$, but instead by the dissipative contribution from the second-order term.
The vanishing of the smallest diagonal shift follows directly from the structure of $V^{D}$. In the vectorized representation,
$V^{D}=H\otimes\mathcal{I}_{R}-\mathcal{I}_{L}\otimes H^{T},$ (29)
so that $V^{D}$ acts as the commutator superoperator, $V^{D}\,|\rho\rangle\rangle=|[H,\rho]\rangle\rangle$. Hence any operator $\rho$ that commutes with $H$ is a zero mode of $V^{D}$. In particular, the identity operator $\mathcal{I}$, any function $f(H)$, and the projectors $|\ket{m}\bra{m}$ in the energy eigenbasis satisfy $H,\rho]=0$, which implies that $V^{D}$ has an exact eigenvalue $\lambda=0$.
[…]
coincide with the corresponding eigenvalues of $V^{D}$ restricted to this manifold, and consequently the smallest value (in magnitude) is
$\min_{k}|\lambda_{k}|=0.$ (31)
It follows that the Liouvillian gap is controlled solely by the second-order term in Eq. (28), yielding the scaling
$\Delta\propto-\,i\,L\,\frac{J^{2}}{\gamma}.$ (32)
@@ Appendix B Appendix B: Reason for choosing the lowest excited state in the rewriting of mixing time
$\tau_{\textit{mix}}(\eta)=\Delta^{-1}\left[\log\left(\operatorname{Tr}|\sigma_{1}|\right)-\log{\left(\frac{2\eta}{|c_{1}|}\right)}\right].$ (33)
@@
A potential concern is that higher excited states, while decaying faster, might have substantially smaller trace norms and could, in principle, yield a larger estimate of the mixing time. For example, one could imagine that for some mode $(\beta_{2},\sigma_{2})$,
$\beta_{2}^{-1}\log\left(\mathrm{Tr}|\sigma_{2}|\right)\;>\;\Delta^{-1}\log\left(\mathrm{Tr}|\sigma_{1}|\right),$ (34)
@@
$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|\left(K_{L}-K_{R}^{*}\right)^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (35)
@@ C.1 Strong-dissipation regime
In the strong-dissipation regime, we assume that $K$ is diagonalizable. The unperturbed doubled-space (vectorized) eigenbasis is $\{|\epsilon_{p},\epsilon_{q}\rangle\}\equiv\{|\epsilon_{p}\rangle_{L}\otimes|\epsilon_{q}\rangle_{R}\}$, where $|\epsilon_{s}\rangle$ are simultaneous eigenstates of the Lindblad operators $\{K_{j}\}$. The leading correction to the state $|\epsilon_{m},\epsilon_{m}\rangle$, induced by the Hamiltonian superoperator
@@
where $H_{pm}\equiv\langle\epsilon_{p}|H|\epsilon_{m}\rangle$, and $\epsilon^{(0)}_{pq}$ denotes the unperturbed doubled-space eigenvalue associated with the basis state $|\epsilon_{p},\epsilon_{q}\rangle$. Here, $\langle\epsilon_{p}|$ denotes the $p$th left eigenvector of $K$. For a general non-Hermitian $K$, the left eigenstate is not necessarily the Hermitian (complex-conjugate) transpose of the corresponding right eigenstate $|\epsilon_{p}\rangle$.
@@ C.2 Weak-dissipation regime
$\begin{split}|E_{m},E_{n}^{(1)}\rangle&=-i\gamma\sum_{p,q}\frac{\langle E_{p},E_{q}^{(0)}|\left(K_{L}-K_{R}^{*}\right)^{2}|E_{m},E_{n}^{(0)}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle\\ &=-i\gamma\sum_{p}\frac{\langle E_{p}|K^{2}|E_{m}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pn}^{(0)}}\,|E_{p},E_{n}^{(0)}\rangle-i\gamma\sum_{q}\frac{\langle E_{q}|K^{2}|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{mq}^{(0)}}\,|E_{m},E_{q}^{(0)}\rangle\\ &\quad+2i\gamma\sum_{p,q}\frac{\langle E_{p}|K|E_{m}\rangle\,\langle E_{q}|K|E_{n}\rangle}{\epsilon_{mn}^{(0)}-\epsilon_{pq}^{(0)}}\,|E_{p},E_{q}^{(0)}\rangle.\end{split}$ (50)
@@
Appendix D Appendix D: Two equivalent ways of rewriting the mixing time
In this appendix, we clarify two equivalent conventions for rewriting the mixing time. The first is the convention adopted in the main text, while the second is commonly used in earlier works. The apparent discrepancy between the two is purely a matter of normalization (i.e., where one “stores” an extensive factor), and does not reflect any physical inconsistency.
One example is the statement in Ref. [60] that the prefactor of the slowest decaying mode scales as $e^{L}$, whereas in our Eq. (11) of the main text the corresponding coefficient $c_{1}$ is at most $O(1)$. These statements are consistent because the product of the coefficient and the mode—which is what enters the physical trace distance—is invariant under rescaling of eigenmodes.
D.1 Way 1 (main text): Hilbert–Schmidt normalization
In the main text we normalize the slowest decaying (right) mode $\sigma_{1}$ by
$\operatorname{Tr}(\sigma_{1}^{2})=1.$ (67)
We then generally write the initial state as
$\rho_{0}=\sigma_{0}+c_{1}\sigma_{1},$ (68)
where $\sigma_{0}$ denotes the steady state.
Because $\rho_{0}$ is a physical density matrix, its purity obeys $\operatorname{Tr}(\rho_{0}^{2})\leq 1$. Expanding and applying Cauchy–Schwarz,
$\operatorname{Tr}(\rho_{0}^{2})$ $=\operatorname{Tr}(\sigma_{0}^{2})+2c_{1}\operatorname{Tr}(\sigma_{0}\sigma_{1})+c_{1}^{2}\operatorname{Tr}(\sigma_{1}^{2})$
$\leq\operatorname{Tr}(\sigma_{0}^{2})+2|c_{1}|\sqrt{\operatorname{Tr}(\sigma_{0}^{2})\operatorname{Tr}(\sigma_{1}^{2})}+|c_{1}|^{2}\operatorname{Tr}(\sigma_{1}^{2})$
$=\operatorname{Tr}(\sigma_{0}^{2})+2|c_{1}|\sqrt{\operatorname{Tr}(\sigma_{0}^{2})}+|c_{1}|^{2}$
$\leq 1+2|c_{1}|+|c_{1}|^{2}=(1+|c_{1}|)^{2}.$ (69)
In the last line we used $\operatorname{Tr}(\sigma_{0}^{2})\leq 1$ and $\sqrt{\operatorname{Tr}(\sigma_{0}^{2})}\leq 1$. Therefore, the physical constraint $\operatorname{Tr}(\rho_{0}^{2})\leq 1$ implies that $|c_{1}|$ cannot grow with system size; in particular, $c_{1}=O(1)$ under our normalization.
With this convention, any potential $e^{L}$ enhancement in the contribution of the slowest mode must come from the trace norm of $\sigma_{1}$, not from $c_{1}$.
D.2 Way 2 (common in the literature): trace-norm normalization
Some references, for example [60, 30] instead normalize the slow mode so that its trace norm is $O(1)$, e.g.
$\operatorname{Tr}|\tilde{\sigma}_{1}|=1,$ (70)
which can be achieved by rescaling
$\tilde{\sigma}_{1}=a\,\sigma_{1},\qquad\tilde{c}_{1}=c_{1}/a.$ (71)
If $\operatorname{Tr}|\sigma_{1}|\sim e^{L}$ in our convention, then choosing $a\sim 1/\operatorname{Tr}|\sigma_{1}|$ makes $\operatorname{Tr}|\tilde{\sigma}_{1}|=O(1)$ and transfers the exponential factor into $\tilde{c}_{1}$. This is exactly how one can obtain an $e^{L}$-scaling coefficient in Ref. [60] while still being consistent with $c_{1}=O(1)$ in our convention.
D.3 Invariant quantity and mixing time
The trace distance to the steady state depends only on the product of the coefficient and the trace norm of the mode. Keeping only the slowest mode for clarity, the time-evolved contribution reads
$\rho(t)-\sigma_{0}=c_{1}\,e^{-i\alpha_{1}t-\Delta t}\,\sigma_{1},$ (72)
and hence
$D(\rho(t),\sigma_{0})=\frac{1}{2}\operatorname{Tr}|\rho(t)-\sigma_{0}|=\frac{|c_{1}|}{2}\,e^{-\Delta t}\,\operatorname{Tr}|\sigma_{1}|.$ (73)
Under the rescaling $\sigma_{1}\mapsto\tilde{\sigma}_{1}=a\sigma_{1}$ and $c_{1}\mapsto\tilde{c}_{1}=c_{1}/a$, the product $|c_{1}|\operatorname{Tr}|\sigma_{1}|$ is unchanged, and therefore the mixing time extracted from (73) is identical in both conventions. In other words, different normalizations merely redistribute the same size dependence between the coefficient and the eigen-operator, without changing any physical prediction.
D.4 Remark on bi-orthogonality of left/right eigenmodes
Finally, we emphasize that in Eq. (11) of the main text, $c_{1}$ is simply the coefficient in the operator decomposition of $\rho_{0}$. We do not identify it with a particular overlap with a left eigenmode, such as
$\tilde{c}_{1}=\langle L_{1}|\rho_{0}\rangle,$ (74)
nor do we impose a normalization condition like $\langle L_{1}|R_{1}\rangle=1$ for left/right eigenmodes of the Lindbladian. Such overlaps and bi-orthogonal normalizations can themselves carry strong system-size dependence for non-normal superoperators. Our bound on $c_{1}$ relies only on the basic physical constraint $\operatorname{Tr}(\rho_{0}^{2})\leq 1$ for a density matrix, together with the Hilbert–Schmidt normalization $\operatorname{Tr}(\sigma_{1}^{2})=1$ used in the main text.